OpenStudy (anonymous):

Find an equation of the tangent line to the given curve at the specified point (3,3) Explain how to do this. y=(5x)/(x+2)

OpenStudy (mimi_x3):

y' = 10/(x+2)^2 (3,3) ; m = 2/5 Tangent: m = 2/5 y-y1 = m(x-x1) y - 3 = 2/5 (x-3) 5y - 15 = 2x - 6 T: 2x - 5y + 9 = 0